OpenStudy (anonymous):

Determine which ordered pairs are part of the solutions set for the inequality. y< x + 2 - A. (-1, 1) B. (0, 3) C. (2, 4) D. (3, 2) Select all that apply

OpenStudy (anonymous):

\[y \le x +2\]

OpenStudy (mathstudent55):

Insert the x and y values of each ordered pair in the inequality, and see which ones make the inequality true. A. (-1, 1) y <= x + 2 Is 1 <= -1 + 2 true? 1 <= 1 True Therefore, (-1, 1) applies Now do the same with the other ordered pairs.

OpenStudy (anonymous):

A and D